# V. Mehrmann and E. MengiRobust Stability of a Quadratic Matrix Polynomial

### v2

@@ 4.2 Interpolation properties
    Suppose $\nu^{(\ell)}$ is a non-degenerate point such that $p_{w}(\|\widehat{z}^{(\ell)}\|)\neq 0$, $g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$ for $\widehat{z}^{(\ell)}:=(\text{Re}(z^{(\ell)}),\text{Im}(z^{(\ell)}))$, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{{\mathcal{V}}_{k}}(\nu^{(\ell)},\widehat{z}^{(\ell)},\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)}))$ is invertible, where $\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)})$ is as in (24) but with $\nu^{(\ell)},\widehat{z}^{(\ell)}$ replacing $\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu})$. Then $\alpha_{\epsilon}(\nu)$, $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ are differentiable at $\nu=\nu^{(\ell)}$, and $\nabla\alpha_{\epsilon}(\nu^{(\ell)})=\nabla\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$.
@@
$\begin{split}\epsilon\;\geq\;\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;&=\;\frac{\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|}{p_{w}(|z^{(\ell)}|)}\\ &\geq\;\frac{\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\,,\end{split}$
@@
$\begin{split}\epsilon\;\geq\;\frac{\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;&=\;\frac{\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|}{p_{w}(|z^{(\ell)}|)}\;\geq\;\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\,,\end{split}$
@@
$\begin{split}\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))\;&=\;\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))\\ \;&=\;\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|\;=\;\|P(z^{(\ell)};\nu^{(\ell)})V_{k}\widetilde{a}\|\,,\end{split}$
@@
As $z^{(\ell)}$ is the unique rightmost point in $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$, the smallest singular value $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ is simple, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{{\mathcal{V}}_{k}}(\nu^{(\ell)},\widehat{z}^{(\ell)},\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)}))$ is invertible by assumption, $\nu^{(\ell)}$ is a non-degenerate point for the restriction of $P(\cdot;\cdot)$ to ${\mathcal{V}}_{k}$. Additionally, by the interpolation of the derivatives in (20), we have $g^{{\mathcal{V}}_{k}}_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})=g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$. Consequently, $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ is differentiable at $\nu=\nu^{(\ell)}$ by Theorem 4.1.
@@ 5 Application to damping optimization
All of the numerical experiments in this section were performed in MATLAB 2024b on a MacBook Air with Mac OS 15.4 operating system, Apple M3 CPU and 16GB RAM using our implementations of Algorithm 3 and 2, as well as other ingredients such as the criss-cross algorithm (i.e., Algorithm 1). The reduced minimization problem in line 3 of Algorithm 3 is solved globally throughout by means of the package EigOpt. When using Algorithm 3, the number of initial interpolation points in line 1 is $\eta=5$ and $\eta=8$ when the problem is dependent on one and two parameters, respectively. The initial interpolation points are chosen as equally spaced points on the interval $\underline{\Omega}$ if there is one parameter or on the line segment that connects the lower-left and upper-right corners of the box $\underline{\Omega}$ if there are two parameters. The prescribed termination tolerance for Algorithm 3 for the examples below is $10^{-8}$.
@@
where $C_{\mathrm{int}}\,=\,\xi 2M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2}$ with $\xi=0.005$.
@@
Computing the globally optimal value with an error less than $10^{-8}$ requires 414 $\epsilon$-pseudospectral abscissa evaluations, and takes about 0.7 seconds. The computed global minimizer $\nu_{\ast}=4.6679$ (i.e., the optimal viscosity) is illustrated in the plots of $\alpha_{\epsilon}(P(\cdot;\nu))$ as a function of the viscosity $\nu$ on the left in Figure 2.
@@
[…]
@@
The iterates $\nu^{(k)}$, $\alpha_{\epsilon}(\nu^{(k)})$ for $\,k\geq 1$ of Algorithm 3 listed in Table 2 indicate a fast convergence. The total runtime of Algorithm 3 until termination is about 1.8 seconds. This is much faster than a direct minimization (i.e., without employing subspace frameworks) of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with Algorithm 1 to compute the objective, which takes about 87.5 seconds. Even a direct minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with the subspace framework (i.e. Algorithm 2) to compute the $\epsilon$-pseudospectral abscissa takes about 25.4 seconds and is still substantially slow. The distribution of the runtime among the different parts of Algorithm 3 is presented in the second, third, fourth columns of Table 5.2. Most of the runtime of Algorithm 3 is taken by the solutions of the reduced problems in line 3, and the computation of the rightmost point in $\Lambda_{\epsilon}(P(\cdot;\nu))$ via Algorithm 2 in lines 1 and 4. Among these two ingredients, solving the reduced problems contributes more to the total runtime of Algorithm 3.
@@
[…]
@@
[…]
@@
$w$ Alg. 3 (with Alg. 2) Alg. 3 (with Alg. 1) direct (with Alg. 2)
$\mathbf{1}$ 12.3 41.4 258.4
$\mathbf{u}$ 18.4 45.9 251.5
[…]
$n$, $\>w$ $\nu_{\ast}$ $\alpha_{\epsilon}(\nu_{\ast})$ $\alpha_{\epsilon}(0)$ iter total reduced psa inner
@@
[…]
@@ 6 Minimizing the pseudospectral abscissa of linearizations
and the corresponding $\epsilon$-pseudospectral abscissa

### v3

@@ 4.2 Interpolation properties
[…]
@@
where $\|\cdot\|$ denotes the Euclidean norm in ${\mathbb{C}}^{\mathsf{n}}$. This shows that $z^{(\ell)}\in\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$, implying that
@@
[…]
@@
[…]
@@
As $z^{(\ell)}$ is the unique rightmost point in $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$, the smallest singular value $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ is simple, and the Hessian ${\mathcal{L}}^{{\mathcal{V}}_{k}}_{yy}(\nu^{(\ell)},\widehat{z}^{(\ell)},\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)}))$ is invertible by assumption, $\nu^{(\ell)}$ is a non-degenerate point for the restriction of $P(\cdot;\cdot)$ to ${\mathcal{V}}_{k}$. Additionally, by the interpolation of the derivatives in (20), we have $g^{{\mathcal{V}}_{k}}_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})=g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$. Consequently, $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ is differentiable at $\nu=\nu^{(\ell)}$ by Theorem 4.1.
@@ 5 Application to damping optimization
All of the numerical experiments in this section were performed in MATLAB 2024b on a MacBook Air with Mac OS 15.4 operating system, Apple M3 CPU and 16GB RAM using our implementations of Algorithm 3 and 2, as well as other ingredients such as the criss-cross algorithm (i.e. Algorithm 1). The reduced minimization problem in line 3 of Algorithm 3 is solved globally throughout by means of the package EigOpt. When using Algorithm 3, the number of initial interpolation points in line 1 is $\eta=5$ and $\eta=8$ when the problem is dependent on one and two parameters, respectively. The initial interpolation points are chosen as equally spaced points on the interval $\underline{\Omega}$ if there is one parameter or on the line segment that connects the lower-left and upper-right corners of the box $\underline{\Omega}$ if there are two parameters. The prescribed termination tolerance for Algorithm 3 for the examples below is $10^{-8}$.
@@
where $C_{\mathrm{int}}\,=\,2\xi M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2}$ with $\xi=0.005$.
@@
Computing the globally optimal value with an error less than $10^{-8}$ requires 414 $\epsilon$-pseudospectral abscissa evaluations, and takes about 0.7 seconds. The computed global minimizer $\nu_{\ast}=4.6679$ (i.e. the optimal viscosity) is illustrated in the plots of $\alpha_{\epsilon}(P(\cdot;\nu))$ as a function of the viscosity $\nu$ on the left in Figure 2.
@@
where as before $C_{\mathrm{int}}\,=\,2\xi M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2}$ with $\xi=0.005$. As in the previous example, we minimize $\alpha_{\epsilon}(P(\cdot;\nu))$ for $\epsilon=0.05$ and $(w_{m},w_{c},w_{k})=(1,1,1)$ over $\nu\in[0,100]$, but now by applying the subspace framework Algorithm 3 with a prescribed error less than $10^{-8}$. After four iterations the method returns $\nu_{\ast}=42.1076$ as the minimizer, as well as $\alpha_{\epsilon}(P(\cdot;\nu_{\ast}))=0.0020$. The accuracy of these computed optimal values is illustrated in the plots on the left-hand side of Figure 3. It is apparent from the plots that, unlike in Example 5.1, the $\epsilon$-pseudospectral abscissa objective is now nonsmooth at the global minimizer, but the method still converges to the global minimizer. The original undamped system is asymptotically stable with $\alpha(P(\cdot;0))=-0.0011$, yet small perturbations move its eigenvalues to the right-half plane, as $\alpha_{\epsilon}(P(\cdot;0))=0.1324$. The optimally damped system has better asymptotic behavior with $\alpha(P(\cdot;\nu_{\ast}))=-0.0079$, and has considerably smaller $\epsilon$-pseudospectral abscissa $\alpha_{\epsilon}(P(\cdot;\nu_{\ast}))=0.0020$, but there are still unstable systems in the $\epsilon$-neighborhood of the optimally damped system. The right-hand plot in Figure 3 depicts the $\epsilon$-pseudospectrum for the undamped and optimally damped system, where again it is evident that optimization moves all of the components of the $\epsilon$-pseudospectrum to the left, but not completely to the left-half of $\mathbb{C}$.
@@
The iterates $\nu^{(k)}$, $\alpha_{\epsilon}(\nu^{(k)})$ for $k\geq 1$ of Algorithm 3 listed in Table 2 indicate a fast convergence. The total runtime of Algorithm 3 until termination is about 1.8 seconds. This is much faster than a direct minimization (i.e. without employing subspace frameworks) of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with Algorithm 1 to compute the objective, which takes about 87.5 seconds. Even a direct minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with the subspace framework (i.e. Algorithm 2) to compute the $\epsilon$-pseudospectral abscissa takes about 25.4 seconds and is still substantially slow. The distribution of the runtime among the different parts of Algorithm 3 is presented in the second, third, fourth columns of Table 5.2. Most of the runtime of Algorithm 3 is taken by the solutions of the reduced problems in line 3, and the computation of the rightmost point in $\Lambda_{\epsilon}(P(\cdot;\nu))$ via Algorithm 2 in lines 1 and 4. Among these two ingredients, solving the reduced problems contributes more to the total runtime of Algorithm 3.
@@
and $C_{\mathrm{int}}\,=\,2\xi M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2},\>\xi=0.005\>$ for n=80,200,400,1200,1400\>n=80,200,400,1200,1400. For every $n$, using Algorithm 3, we minimize $\alpha_{\epsilon}(\nu)$ over $\nu\in[0,250]$ for $\epsilon=0.03$ and two different choices for the weights, namely $(w_{m},w_{c},w_{k})=(1,1,1)$ and $(w_{m},w_{c},w_{k})=(0.7,1,0)$. Note that for the latter choice, perturbations of the stiffness matrix are not allowed. Physically, the parameter-dependent matrix polynomial again corresponds to a mass-spring-damper system consisting of $n$ masses connected with springs such that the second and third mass, as well as the fourth and fifth mass are tied through dampers. Moreover, both of the two dampers are required to have the same viscosity.
@@
As we have observed before, also looking at the times reported in the columns of ‘total’, ‘reduced’, ‘psa’ of Table 5, the total runtime is more or less determined by (i) the time to solve the reduced problems and (ii) the time to locate the rightmost points in the $\epsilon$-pseudospectrum of the full problems. In these test examples for smaller matrix polynomials, the former contributes significantly more to the computation time. However, the contribution of the latter becomes substantial as the sizes of the matrices increase. For larger matrix polynomials with $n>1000$, according to Table 5, computing the rightmost points in the $\epsilon$-pseudospectrum of the full problems via Algorithm 2 dominates the runtime. In such a call to Algorithm 2 for the computation of $\alpha_{\epsilon}(P(\cdot;\nu))$ $\exists\nu\in\underline{\Omega}\,$ and the parameter-dependent matrix polynomial $P$ with $n=1200$ or $n=1400$, at least about $\%95$ of the total time is taken by the large-scale eigenvalue and singular value computations in lines 1, 2, 6 of Algorithm 2.
@@
$n$, w\>w $\nu_{\ast}$ $\alpha_{\epsilon}(\nu_{\ast})$ $\alpha_{\epsilon}(0)$ iter total reduced psa inner
@@
$w$ Alg. 3 (with Alg. 2) Alg. 3 (with Alg. 1) direct (with Alg. 2)
$\mathbf{1}$ 12.3 41.4 258.4
$\mathbf{u}$ 18.4 45.9 251.5
Table 6: Runtimes in seconds to minimize $\alpha_{\epsilon}(P(\cdot;\nu))$ for the damping problem with $n=80$ in Example 5.3 and the weights $w=(w_{m},w_{c},w_{k})$ equal to either ${\mathbf{1}}=(1,1,1)$ or ${\mathbf{u}}=(0.7,1,0)$.
@@
with $C_{\mathrm{int}}(\nu_{2})\,=\,\,2\xi M^{1/2}\sqrt{M^{-1/2}K(\nu_{2})M^{-1/2}}M^{1/2},\>\xi=0.005\>$. Thus, now there is only one damper on the second mass, whose viscosity $\nu_{1}$ is an optimization parameter. Additionally, the spring constant $\nu_{2}$ affecting the stiffness matrix is the second optimization parameter. We minimize $\alpha_{\epsilon}(\nu)$ for $\epsilon=0.05$ and $(w_{m},w_{c},w_{k})=(1,1,1)$ over $\nu=(\nu_{1},\nu_{2})\>\in\>[0,100]\times[20,80]$ using Algorithm 3. The computed global minimizer is $\nu_{\ast}=(72.4622,80)$. The right-hand plot in Figure 4 illustrates the accuracy of this computed minimizer. Additionally, quantities related to the application of Algorithm 3 are provided in Table 8. The computed globally smallest value of the $\epsilon$-pseudospectral abscissa is now $\alpha_{\epsilon}(\nu_{\ast})=-0.0081$, and again indicates uniform asymptotic stability in an $\epsilon$ neighborhood of the optimal system, unlike the minimal value $0.0020$ when minimization is performed only with respect to the damping parameter on the second mass.
@@ 6 Minimizing the pseudospectral abscissa of linearizations
with $\Lambda({\mathcal{A}}+\Delta)$ denoting the set of eigenvalue of the matrix ${\mathcal{A}}+\Delta$, and the corresponding $\epsilon$-pseudospectral abscissa